OpenStudy (anonymous):

2-3(5-x)=3x+1 I need to solve for x, but does this equation have no solution?

OpenStudy (anonymous):

No solution.

OpenStudy (anonymous):

diddo

OpenStudy (anonymous):

no solution

OpenStudy (anonymous):

no wait it would be no real solution

OpenStudy (amistre64):

2-3(5-x)=3x+1 -2 -2 ---------------- -3(5-x) = 3x -1 /-3 /-3 ------------- 5-x = -x +1/3 +x +x ------------- 5 = 1/3 ; cant be true ever

OpenStudy (amistre64):

by real standards ;)